Commercial electrolytic cells for producing aluminum



operate at 5.0 V and 100,000 A.



(a) How long does it take to produce exactly 1 metric ton (1000 kg)



of aluminum?



(b) How much electrical power (in kilowatt-hours, kW·h) is used



(1 W = 1 J/s; 1 kW·h = 3.6×103



kJ)?



(c) If electricity costs $0.123 per kW·h and cell efficiency is 90.%,



what is the cost of electricity to produce exactly 1 lb of aluminum?

1
Expert's answer
2021-11-22T17:01:03-0500

(a) The half cell reaction is:Al3+ + 3e- ______ Al

Mass of Al deposited =1000kg×1000g/kg

According to the balanced equation 3 mol of electrons involved in producing one mole of Al i.e. 27 g

3 mol - 27 g 

? - 

Moles of e- transferred

= 1.11×105mol

Charge of 1 mol of  = 96500

Charge (C) = 1.11×105

96500 ×11100=1.07×10115

Time =Quanrity /f

Time=1.073×105s

B)power =I×v

=100000×5

=500kw